# Leveraging Sidewalk Robots for Walkability-Related Analyses

### v1

@@ Abstract
Walkability is a key component of sustainable urban development, while collecting detailed data on its related features remains challenging due to the high costs and limited scalability of traditional methods. Sidewalk delivery robots, increasingly deployed in urban environments, offer a promising solution to these limitations. This paper explores how these robots can serve as mobile data collection platforms, capturing sidewalk-level features related to walkability in a scalable, automated, and real-time manner. A sensor-equipped robot was deployed on a sidewalk network at KTH in Stockholm, completing 101 trips covering 900 segments. From the collected data, different typologies of features are derived, including robot trip characteristics (e.g., speed, duration), sidewalk conditions (e.g., width, surface unevenness), and sidewalk utilization (e.g., pedestrian density). Their walkability-related implications were investigated with a series of analyses. The results demonstrate that pedestrian movement patterns are strongly influenced by sidewalk characteristics, with higher density, reduced width, and surface irregularity associated with slower and more variable trajectories. Notably, robot speed closely mirrors pedestrian behavior, highlighting its potential as a proxy for assessing pedestrian dynamics. The proposed framework enables continuous monitoring of sidewalk conditions and pedestrian behavior, contributing to the development of more walkable, inclusive, and responsive urban environments.
@@ 1 Introduction
Whether operational or perceived, walkability is increasingly recognized as a vital aspect of sustainable urban development. To develop more walkable cities, one must first unravel the complex pedestrian mobility patterns and interactions with existing infrastructure (Rhoads et al.,, 2023). Thus, increasing efforts are being made in assessing the usage of pedestrian infrastructure, identifying critical features (e.g., pedestrian density, pedestrian speed), and understanding the complex pedestrian mobility patterns and interactions (Fruin,, 1970; Maghelal and Capp,, 2011; Dragović et al.,, 2023). Currently, pedestrian flow data is typically collected through infrared (IR) break-beam sensors and cameras placed at specific locations across the city (Moussaïd et al.,, 2010; Alahi et al.,, 2014; Carter et al.,, 2020). While this type of solution can offer granular data, it is typically difficult to adopt it on a broad scale due to its high costs and privacy issues (e.g., General Data Protection Regulation in Europe). Other studies focusing on the larger-scale analysis of urban mobility patterns (e.g., origins and destinations, mode choice, trip features) rely on GPS, wireless and cellular networks, and social media geotag records (Leslie et al.,, 2007; Kung et al.,, 2014; D’Orso and Migliore,, 2020; Gao et al.,, 2014). These solutions offer a good overview of pedestrian mobility at an aggregate level, but they cannot capture more complex phenomena such as, pedestrian flows, trajectories, and interactions with other road users. Surveys and audits are also used for collecting qualitative data on perceived walkability (Steve and Shane,, 2011; Cambra and Moura,, 2020; Chen et al.,, 2024) and quantitative measures of the built environment (Bartzokas-Tsiompras et al.,, 2021; Lee et al.,, 2020). Although these methods provide useful insights, they are time-consuming, resource-intensive, and difficult to scale, posing a major bottleneck to advancing the knowledge of walkability in research.
@@
Building on this trend of leveraging emerging technologies for walkability analysis, sidewalk delivery robots offer a novel and scalable opportunity to address key data limitations by enabling continuous, real-time collection of sidewalk-level information. Equipped with a suite of onboard sensors and real-time processing capabilities, these robots can autonomously collect and analyze data during routine deliveries. In this study, we illustrate which sidewalk walkability-related features can be collected. These features are categorized into three primary domains: robot trip features (e.g., robot speed, travel duration), sidewalk condition features (e.g., effective walkway width, unevenness), and sidewalk utilization features (e.g., pedestrian density and speed). Walkability is inherently multifaceted, shaped by social, cultural, and individual preferences. Although sidewalk robots cannot capture all factors influencing walkability, they can provide valuable insights into the walking infrastructure quality, spatio-temporal pedestrian patterns, and measurable aspects of the walking experience.
This study does not aim to define walkability, but rather identifies measurable walkability-related factors through automatic robot-based data collection and uses them for pedestrian mobility assessment. As sidewalk delivery robots become more common in urban environments, their routine movements offer a practical means to evaluate pedestrian mobility dynamically and at scale. Compared to traditional methods, this approach is significantly more time- and resource-efficient, enabling continuous and automated assessments wherever robots can travel.
The remainder of this paper is structured as follows. In Section 2, we briefly review the relevant literature on features that influence sidewalk walkability. Section 3 describes the data collection procedures and the methodology used to derive walkability-related features from raw robot sensor data. Section 4 includes several analyses of the extracted features, exploring differences in sidewalk conditions and their utilization, pedestrian behavior patterns, and the influence of sidewalk characteristics on pedestrian movement. Finally, Section 5 summarizes the main findings and discusses potential directions for future research.
@@ 2 Related Work
While these studies offer a broad conceptual and empirical understanding of walkability, the present study narrows the focus to operational features, which refer to specific environmental factors that directly affect the experience of walking from a functional perspective. Based on a review of previous studies, we have identified twelve relevant factors that can potentially affect walkability from an operational perspective: 1) static obstacles; 2) dynamic obstacles; 3) walkway width; 4) walkway quality; 5) sidewalk slope; 6) crosswalk distance; 7) crosswalk type; 8) potential for conflicts; 9) vehicle speed; 10) traffic volume; 11) driveways; 12) supporting facilities. Among these, factors 6 to 12 are mainly related to safety and crossing conditions. Each feature may correspond to multiple indicators or measurements across different studies.
Obstacles in the walkway can cause deviation from the original path, which is proven to be one of the major factors that affect the walkability of a path according to Steve and Shane, (2011). Reisi et al., (2019) applied the Analytical Hierarchy Process (AHP) for weighting indicators of walkability. The number of static obstacles (e.g. poles, signs, chairs along sidewalks) was found to have high relative weight. Other static obstacles, such as wayfinding, green hedge, trees, trash cans, public seating furniture, and illegally parked vehicles, were also considered to have an influence on walkability (Yin,, 2017; Mitropoulos et al.,, 2023; Bartzokas-Tsiompras et al.,, 2021). In addition to using the number of obstacles as a measure, Golan et al., (2019) employed the number of requests for cleanup of streets, garbage cans, bulky items, and other waste, which was ranked as the third most influential variable on women’s walkability.
Aside from static obstacles, the impact of dynamic obstacles on walkability, especially for pedestrians, should not be underestimated on sidewalks. Fruin, (1970) and D’Orso and Migliore, (2020) used pedestrian flow to describe the pedestrian level of service (LOS), which is an indicator of walkability. Arellana et al., (2020) incorporated pedestrian volume as one of the components in assessing the safety factor when measuring walkability. The observed variation in pedestrian volume can also be utilized to indicate or symbolize the alteration in walkability (Cambra and Moura,, 2020). Moreover, the results of Jardim et al., (2023) revealed that the strength of the relationship between pedestrian flow and walkability differs between weekdays and weekends.
@@
The walkability improves as the physical road conditions enhance (Steve and Shane,, 2011), thus the quality of the walkway can be qualitatively evaluated based on regularity, smoothness, and slippery characteristics of the pavement (Corno and Savaresi,, 2020; Moura et al.,, 2017), and is often evaluated through field audits (Bartzokas-Tsiompras et al.,, 2021; Lee et al.,, 2020). Closely related to surface quality is the sidewalk slope, which affects both walking comfort and accessibility, especially for individuals with mobility limitations. Slope can be measured in two primary ways: by calculating the average gradient along a segment (Deng et al.,, 2020; Taleai and Yameqani,, 2018), or by measuring the elevation range between the highest and lowest points on the segment (Fan et al.,, 2018).
@@
Supporting facilities encompass elements such as ramps or steps at the intersection of sidewalks and crosswalks, central islands within crosswalks, and sidewalk buffers. The presence of ramps or steps benefits individuals with disabilities (Mulyadi et al.,, 2022). Additionally, the presence of central islands and buffers is crucial for ensuring walkway safety, which in turn contributes to overall walkability (Bracy et al.,, 2014; Bartzokas-Tsiompras et al.,, 2021; Steve and Shane,, 2011).
@@ 3 Walkability-Related Features
This section outlines the data collection process and the methodology used to derive walkability-related features from raw robot sensor data. We first describe the experimental setup, including the sidewalk robot platform and its deployment across the sidewalk network. The subsequent subsections present three key categories of features extracted from the collected data: robot trip features that characterize movement behavior, sidewalk condition features that reflect surface quality and geometry, and sidewalk utilization features that capture patterns of pedestrian interaction with the environment.
@@ 3.1 Experimental setup
The robot utilized for sidewalk data collection is a custom-designed, open-source vehicle called ”Small Vehicles for Autonomy” (SVEA) (Jiang et al.,, 2022), as shown in Figure 1(a). Although the testbed of SVEA is designed to support automated driving systems under V2X communication, in this study, the robot was manually driven along sidewalks to collect data for safety reasons. The platform is built on the Traxxas TRX-4, a four-wheel-drive vehicle equipped with a two-gear transmission system. To operate the robot in a manner similar to a typical sidewalk autonomous delivery robot, which travels at pedestrian speeds, we selected the low-gear mode and programmatically restricted its velocity during the data collection experiment. With these settings, the maximum speed of the robot in low gear was capped at approximately 1.6 m/s. The robot’s footprint measures 28 cm × 57 cm, and its weight is 6.5 kg. It is equipped with a suite of onboard sensors, including a Stereolabs ZED 1 stereo camera, a u-blox ZED-F9P GNSS module for RTK-GPS, a Hokuyo UST-10LX LiDAR, and an off-the-shelf Adafruit IMU, to achieve real-time perception and localization. On-board computing is facilitated by the Stereolabs ZED-BOX, which contains a Jetson Xavier NX, serving as the central computing unit that processes sensor data in real time.
@@
(a) SVEA robot
(b) Sidewalk network
Figure 1: The SVEA robot used for sidewalk data collection (a) and the mini network on KTH campus where the robot operates (b)
The robot was deployed on a small sidewalk network within the KTH campus in Stockholm, Sweden, as shown in Figure 1(b). The network consists of multiple segments, including both sidewalks and pedestrian crossings, designed to capture a variety of walking environments. Each segment serves as a unit of analysis for evaluating sidewalk features across the area.
From September 18, 2024, to March 24, 2025, the robot completed 101 trips, covering 900 records of sidewalk segments in total. Trips were conducted on different days and at various times between 08:00 and 18:00, with particular focus on peak hours. During each trip, onboard sensors continuously recorded raw data, including GNSS points, robot velocity and heading, IMU readings, sidewalk width, ambient brightness, and anonymized pedestrian data. This dataset was processed to derive three categories of walkability-related features: robot trip features, sidewalk condition features, and sidewalk utilization features. These extracted features offer valuable insights into the quality of walking infrastructure and spatio-temporal pedestrian patterns. An overview of these features is presented in Table 1.

### v2

@@ Abstract
Walkability is a key component of sustainable urban development, while collecting detailed data on sidewalks (or pedestrian infrastructures) remains challenging due to the high costs and limited scalability of traditional methods. Sidewalk delivery robots, increasingly deployed in urban environments, offer a promising solution to these limitations. This paper explores how these robots can serve as mobile data collection platforms, capturing sidewalk-level features related to walkability in a scalable, automated, and real-time manner. A sensor-equipped robot was deployed on a sidewalk network at KTH in Stockholm, completing 101 trips covering 900 segment records. From the collected data, different typologies of features are derived, including robot trip characteristics (e.g., speed, duration), sidewalk conditions (e.g., width, surface unevenness), and sidewalk utilization (e.g., pedestrian density). Their walkability-related implications were investigated with a series of analyses. The results demonstrate that pedestrian movement patterns are strongly influenced by sidewalk characteristics, with higher density, reduced width, and surface irregularity associated with slower and more variable trajectories. Notably, robot speed closely mirrors pedestrian behavior, highlighting its potential as a proxy for assessing pedestrian dynamics. The proposed framework enables continuous monitoring of sidewalk conditions and pedestrian behavior, contributing to the development of more walkable, inclusive, and responsive urban environments.
@@ 1 Introduction
Whether operational or perceived, walkability is increasingly recognized as a vital aspect of sustainable urban development. To develop more walkable cities, one must first unravel the complex pedestrian mobility patterns and interactions with existing infrastructure (Rhoads et al.,, 2023). Thus, increasing efforts are being made in assessing the usage of pedestrian infrastructure, identifying critical features (e.g., pedestrian density, pedestrian speed), and understanding the complex pedestrian mobility patterns and interactions (Fruin,, 1970; Maghelal and Capp,, 2011; Dragović et al.,, 2023). Currently, pedestrian flow data is typically collected through infrared (IR) break-beam sensors and cameras placed at specific locations across the city (Moussaïd et al.,, 2010; Alahi et al.,, 2014; Carter et al.,, 2020). While this type of solution can offer granular data, it is typically difficult to adopt it on a broad scale due to its high costs and privacy issues (e.g., General Data Protection Regulation in Europe). More recently, street-level imagery (e.g., Google Street View) has been explored as an alternative source, where pedestrians can be detected and counted with computer vision techniques (Liu et al.,, 2023; Chen et al.,, 2020). These imagery-based approaches are less costly to deploy and scalable across large areas, but they suffer from limited temporal resolution, since imagery is updated infrequently and cannot provide continuous measurements. Other studies focusing on the larger-scale analysis of urban mobility patterns (e.g., origins and destinations, mode choice, trip features) rely on GPS, wireless and cellular networks, and social media geotag records (Leslie et al.,, 2007; Kung et al.,, 2014; D’Orso and Migliore,, 2020; Gao et al.,, 2014). These solutions offer a good overview of pedestrian mobility at an aggregate level, but they cannot capture more complex phenomena such as, pedestrian flows, trajectories, and interactions with other road users. Surveys and audits are also used for collecting qualitative data on perceived walkability (Steve and Shane,, 2011; Cambra and Moura,, 2020; Chen et al.,, 2024) and quantitative measures of the built environment to develop walkability indicators (Bartzokas-Tsiompras et al.,, 2021; Lee et al.,, 2020). Although these methods provide useful insights, they are time-consuming, resource-intensive, and difficult to scale, posing a major bottleneck to advancing the knowledge of walkability in research.
@@
Building on this trend of leveraging emerging technologies for walkability analysis, sidewalk delivery robots, which are currently among the most common real-world applications of autonomous ground vehicles, offer a novel and scalable opportunity to address key data limitations by enabling continuous, real-time collection of sidewalk-level information. Recent studies have explored these robots from multiple perspectives, including their interactions and potential conflicts with pedestrians (Gehrke et al.,, 2023), the regulatory challenges associated with mitigating disruptions and risks (Kovacic et al.,, 2024), and their role in meeting the growing demand for local deliveries and e-commerce (Yang et al.,, 2025). Equipped with a suite of onboard sensors and real-time processing capabilities, these robots can autonomously collect and analyze data during routine deliveries. In this study, we illustrate how raw sensor data can be transformed into sidewalk walkability-related features at high spatial and temporal resolution. These features are categorized into three primary domains: robot trip features (e.g., robot speed, travel duration), sidewalk condition features (e.g., effective walkway width, unevenness), and sidewalk utilization features (e.g., pedestrian density and speed). Walkability is inherently multifaceted, shaped by social, cultural, and individual preferences. Although sidewalk robots cannot capture all factors influencing walkability, they can provide valuable insights into the walking infrastructure quality, spatio-temporal pedestrian patterns, and measurable aspects of the walking experience.
This study does not aim to define walkability. Instead, it demonstrates how sidewalk robots can be leveraged to automatically capture measurable walkability-related features and use them to assess pedestrian mobility in practice. The contributions of this study are threefold. First, it introduces robots as a novel data collection instrument, capable of overcoming limitations of traditional methods which prevent their application and scalability. As sidewalk delivery robots become more common in urban environments, their routine movements offer a practical and cost-efficient means to evaluate pedestrian mobility real-time. Second, by grouping different operational, walkability-related factors into robot trip features, sidewalk condition features, and sidewalk utilization features, the study establishes a systematic framework for extracting and analyzing data. Third, it provides empirical evidence that robot-based measurements can both confirm established relationships in pedestrian dynamics and reveal context-specific patterns, such as how width, density, and surface quality shape walking speed and movement styles. In the long run, this robot-based method can help build richer and more inclusive datasets that complement traditional audits and fixed sensors.
The remainder of this paper is structured as follows. In Section 2, we briefly review the relevant literature on features that influence sidewalk walkability. Section 3 describes the data collection procedures and the methodology used to derive walkability-related features from raw robot sensor data. Section 4 includes several analyses of the extracted features, exploring differences in sidewalk conditions and their utilization, pedestrian behavior patterns, and the influence of sidewalk characteristics on pedestrian movement. Section 5 offers deep reflections on our findings and discusses potential directions for future research. Finally, Section 6 summarizes the main findings, highlights contributions. and outlines limitations and future research directions.
@@ 2 Related Work
While these studies offer a broad conceptual and empirical understanding of walkability, this study focuses on the operational features, specific environmental factors that directly affect the functional experience of walking. Rather than a systematic survey of the entire walkability literature, we summarize relevant across complementary fields, including urban planning, transport geography, pedestrian mobility studies, and traffic flow theory. The selected studies explicitly examined the relation between the built environment and pedestrian movements, provided measurable indicators at the sidewalk level, and offered empirical insights. On this basis , we identified twelve relevant factors from an operational perspective: 1) static obstacles; 2) dynamic obstacles; 3) walkway width; 4) walkway quality; 5) sidewalk slope; 6) crosswalk distance; 7) crosswalk type; 8) potential for conflicts; 9) vehicle speed; 10) traffic volume; 11) driveways; 12) supporting facilities. Among these, factors 6 to 12 are mainly related to safety and crossing conditions. Each feature may correspond to multiple indicators or measurements across different studies.
Obstacles in the walkway can cause deviation from the original path, which is proven to be one of the major factors that affect the walkability of a path according to Steve and Shane, (2011). Reisi et al., (2019) applied the Analytical Hierarchy Process (AHP) for weighting indicators of walkability. The number of static obstacles (e.g. poles, signs, chairs along sidewalks) was found to have high relative weight. Other static obstacles, such as wayfinding, green hedge, trees, trash cans, public seating furniture, and illegally parked vehicles, were also considered to have an influence on walkability (Yin,, 2017; Mitropoulos et al.,, 2023; Bartzokas-Tsiompras et al.,, 2021). In addition to using the number of obstacles as a measure, Golan et al., (2019) employed the number of requests for cleanup of streets, garbage cans, bulky items, and other waste, which was ranked as the third most influential variable on women’s walkability. Beyond the obstacles’ general impact on route continuity and convenience, they can pose significant hazards for people with disabilities, particularly blind or visually impaired pedestrians.
Aside from static obstacles, the impact of dynamic obstacles on walkability should not be underestimated. Dynamic obstacles refer to temporary or moving obstructions that interrupt or constrain the walking experience. These may include illegally parked cars, delivery vehicles making curb stops, scattered e-scooters and bicycles, street vendors or construction equipment occupying sidewalks, and, most commonly, other pedestrians. Several studies have examined how pedestrian presence affects walkability. Fruin, (1970) and D’Orso and Migliore, (2020) used pedestrian flow to describe the pedestrian level of service (LOS), which is an indicator of walkability. Arellana et al., (2020) incorporated pedestrian volume as one of the components in assessing the safety factor when measuring walkability. The observed variation in pedestrian volume can also be utilized to indicate or symbolize the alteration in walkability (Cambra and Moura,, 2020). Moreover, the results of Jardim et al., (2023) revealed that the strength of the relationship between pedestrian flow and walkability differs between weekdays and weekends.
@@
The walkability improves as the physical road conditions enhance (Steve and Shane,, 2011), thus the quality of the walkway can be qualitatively evaluated based on regularity, smoothness, and slippery characteristics of the pavement (Corno and Savaresi,, 2020; Moura et al.,, 2017), and is often evaluated through field audits (Bartzokas-Tsiompras et al.,, 2021; Lee et al.,, 2020). In addition to these attributes, pavement friction and texture are also important characteristics influencing pedestrian movement. Together, these surface properties not only shape comfort and accessibility but also play a critical role in safety, as poor conditions such as uneven or slippery pavements can increase the risk of pedestrian falls (Silva et al.,, 2025). Closely related to surface quality is the sidewalk slope, which affects both walking comfort and accessibility, especially for individuals with mobility limitations. Slope can be measured in two primary ways: by calculating the average gradient along a segment (Deng et al.,, 2020; Taleai and Yameqani,, 2018), or by measuring the elevation range between the highest and lowest points on the segment (Fan et al.,, 2018).
@@
Supporting facilities encompass elements such as ramps or steps at the intersection of sidewalks and crosswalks, central islands within crosswalks, and sidewalk buffers. Ramps facilitate accessibility and are particularly beneficial for individuals with disabilities (Mulyadi et al.,, 2022), whereas steps are often considered barriers, especially for older adults, wheelchair users, or parents with strollers. Additionally, the presence of central islands and buffers is crucial for ensuring walkway safety, which in turn contributes to overall walkability (Bracy et al.,, 2014; Bartzokas-Tsiompras et al.,, 2021; Steve and Shane,, 2011).
@@ 3 Walkability-Related Features
This section outlines the data collection process and the methodology used to derive walkability-related features from raw robot sensor data. We first describe the experimental setup, including the sidewalk robot platform and its deployment across the sidewalk network. The subsequent subsections present three key categories of features extracted from the collected data: robot trip features that characterize movement behavior, sidewalk condition features that reflect surface quality and geometry, and sidewalk utilization features that capture patterns of pedestrian interaction with the environment. Figure 1 provides an overview of the data collection workflow and processing pipeline, illustrating how these components contribute to the walkability analysis.
Figure 1: Work flow of sidewalk robot-based data collection and processing for walkability analysis
@@ 3.1 Experimental setup
The robot utilized for sidewalk data collection is a custom-designed, open-source vehicle called ”Small Vehicles for Autonomy” (SVEA) (Jiang et al.,, 2022), as shown in Figure 2. Although the testbed of SVEA is designed to support automated driving systems under V2X communication, in this study, the robot was manually driven along sidewalks to collect data for safety reasons. The platform is built on the Traxxas TRX-4, a four-wheel-drive vehicle equipped with a two-gear transmission system. To operate the robot in a manner similar to a typical sidewalk autonomous delivery robot, which travels at pedestrian speeds, we selected the low-gear mode and programmatically restricted its velocity during the data collection experiment. With these settings, the maximum speed of the robot in low gear was capped at approximately 1.6 m/s. The robot’s footprint measures 28 cm × 57 cm, and its weight is 6.5 kg. It is equipped with a suite of onboard sensors, including a Stereolabs ZED 1 stereo camera, a u-blox ZED-F9P GNSS module for RTK-GPS, a Hokuyo UST-10LX LiDAR, and an off-the-shelf Adafruit IMU, to achieve real-time perception and localization. On-board computing is facilitated by the Stereolabs ZED-BOX, which contains a Jetson Xavier NX, serving as the central computing unit that processes sensor data in real time.
@@
Figure 2: SVEA robot
(a) Sidewalk network
(b) Network location within Stockholm
Figure 3: Network used in the experiments on KTH campus
The robot was deployed on a small sidewalk network within the KTH campus in Stockholm, Sweden, as shown in Figure 3. The network consists of multiple segments, including both sidewalks and pedestrian crossings, designed to capture a variety of walking environments. For each street in the selected area, both sidewalks were included in the analysis rather than limiting the study to a single side. As illustrated in Figure 3(a), the sidewalks on the southwest side have more points of interest compared to those on the northeast side, resulting in different levels of pedestrian activity across the network. Each segment therefore serves as a unit of analysis for evaluating sidewalk features across the area.
KTH campus was selected as the study area due to several reasons. As an urban campus, it offers a diverse range of pedestrian facilities, and stable pedestrian volumes. The selected area, which exhibits high network centrality within the campus sidewalk system, is traversed daily by most students and staff, making it an high-activity pathway. Its combination of academic buildings, green spaces, and service facilities generates periodic pedestrian flows. These characteristics create dynamic settings that support realistic testing, while also offering a manageable environment for repeated robot deployments. Safety and practicality also played a role since the robot needed to be supervised during the experiments. A concise overview of the network characteristics is provided in Table 1 to support interpretation and transferability.
Table 1: Summary of sidewalk network characteristics in the KTH campus study area.
Category Description
Scale and composition Small-scale network including both sidewalks for each street, and two pedestrian crossings.
Activity and usage High centrality within campus; daily flows of students and staff; periodic peaks during class hours; southwest side with higher activity due to points of interest.
From September 18, 2024, to March 24, 2025, the robot completed 101 trips, covering 900 records of sidewalk segments in total. Trips were conducted on different days and at various times between 08:00 and 18:00, with particular focus on peak hours. During each trip, onboard sensors continuously recorded raw data, including GNSS points, robot velocity and heading, IMU readings, sidewalk width, ambient brightness, and anonymized pedestrian data. The experiments were conducting during months of higher pedestrian activity on campus, ensuring the dataset’s relevance. This dataset was processed to derive three categories of walkability-related features: robot trip features, sidewalk condition features, and sidewalk utilization features. These extracted features offer valuable insights into the quality of walking infrastructure and spatio-temporal pedestrian patterns. An overview of these features is presented in Table 2, while an interactive dashboard [^1] https://ismir-sidewalk-mobility-data-dashboard.streamlit.app/ visualizes sidewalk usage and characteristics for the study area.